# Improved Convergence Rates of Muon Optimizer for Nonconvex Optimization

### v1

@@ 1.1 Background
Muon (Jordan et al., 2024) performs updates based on first-order information, similarly to stochastic gradient descent (SGD) (Robbins & Monro, 1951) and its momentum variants (Polyak, 1964; Nesterov, 1983; Sutskever et al., 2013), while distinguishing itself by incorporating an orthogonalization operation into the update direction (see (Si et al., 2025; Zhang et al., 2025; Huang et al., 2025; Liu et al., 2025b; Gruntkowska et al., 2025; Ahn et al., 2025) for the variants of Muon). Whereas Adam (Kingma & Ba, 2015) stabilizes training through coordinate-wise adaptive scaling, Muon projects the update matrix onto its orthogonal factor, thereby explicitly controlling the correlation structure of the update direction. This property enables Muon to preserve directional information even in high-dimensional spaces, achieving both numerical stability and efficient optimization. Empirical studies have further reported that Muon has advantages over Adam-type optimizers (Kingma & Ba, 2015; Loshchilov & Hutter, 2019) in terms of both computational efficiency and scalability (Liu et al., 2025a; Wang et al., 2025; AI et al., 2025; Mehta et al., 2025), highlighting its potential as a new optimizer capable of replacing conventional methods.
@@ 1.2 Motivation
Several attempts (Tang et al., 2025; Chang et al., 2025; Sato et al., 2025; Shen et al., 2025; Li & Hong, 2025) have been made to theoretically analyze the convergence properties of the Muon optimizer; the existing results are summarized in Table 1. These analyses aim to clarify the theoretical characteristics of Muon by establishing convergence guarantees or upper bounds under specific assumptions. However, each of these analyses is subject to limitations stemming from the strength of the imposed assumptions or the choice of evaluation criteria, and thus it does not necessarily fully capture the actual update structure of Muon.
Table 1: Comparison of the existing results (1)–(5) with our results (R1)(R5) for Muon. The existing results are (1) (Tang et al., 2025, Theorem 4.6), (2) (Chang et al., 2025, Theorem 3.4), (3) (Sato et al., 2025, Theorem 3.1), (4) (Shen et al., 2025, Theorem 4.1), and (5) (Li & Hong, 2025, Theorem 2.1), where Result (2) assumes the PL condition. $(\bm{W}_{t})$ is the sequence generated by Muon. $\|\cdot\|_{\mathrm{F}}$ is for the Frobenius norm and $\|\cdot\|_{*}$ is the dual norm of $\|\cdot\|_{\mathrm{F}}$. $\mathbb{E}$ is the total expectation. Let $\eta>0$, $b>0$, $\delta>1$, and $T$ be the number of steps. $O$ stands for Landau’s symbol, i.e., $y_{T}=O(x_{T})$ implies that, there exist $C>0$ and $t_{0}\in\mathbb{N}$ such that, for all $T\geq t_{0}$, $y_{T}\leq Cx_{T}$, while $y_{T}=\mathcal{O}(x_{T})$ implies that, for a fixed $T$, $y_{T}\leq Cx_{T}$. The notation $O(x(T,\eta,b))\hookrightarrow\mathcal{O}(x(T))$ implies that, when $\eta$ and $b$ are set by using the fixed $T$, the convergence rate changes from $O(x(T,\eta,b))$ to $\mathcal{O}(x(T))$.
@@
Upon examining the results summarized in Table 1, we observe that Result (1) (Tang et al., 2025, Theorem 4.6) provides the convergence guarantee of Muon; however, its convergence rate would be inferior to that obtained in the present work. Result (2) (Chang et al., 2025, Theorem 3.4) can be interpreted as achieving a good convergence rate $O(\frac{(\log{T})^{2}}{T^{2/3}})$; however, it relies on the strong Polyak–Łojasiewicz (PL) condition, which significantly restricts its range of applicability. In the case of (3) (Sato et al., 2025, Theorem 3.1), a non-negligible variable $n$, where a parameter of a DNN is in $\mathbb{R}^{m\times n}$, remain in the analysis, and therefore a complete convergence result in a strict sense is not established. Result (4) (Shen et al., 2025, Theorem 4.1), similar to (2), provides a certain convergence guarantee, but its convergence rate would be again inferior to that achieved in this work. Result (5) (Li & Hong, 2025, Theorem 2.1) indicates that setting the number of steps $T$ as the batch size $b$ ensures a better $\mathcal{O}(\frac{1}{\sqrt{T}})$ convergence rate of Muon than an $\mathcal{O}(\frac{1}{T^{1/4}})$ convergence rate in Results (1) and (4).
@@ 1.3 Contribution
Improved convergence rate for the Muon optimizer in the sense of the notation $\mathcal{O}$: We evaluate the upper bound in the sense of the notation $\mathcal{O}$, i.e., the case where the number of steps $T$ needed to train the DNN is fixed, by using practical learning rates and batch sizes. For example, when using a learning rate $\eta_{t}$ $(\leq\eta)$ defined by a constant, cosine-annealing, or polynomial decay learning rate and a constant batch size $b_{t}=b$, Muon has the upper bound $O(\frac{1}{T}+\eta+\frac{1}{\sqrt{b}})$, as shown in (R1) and (R2) in Table 1 (Corollary 3.1(i), (iii), and (v)). Hence, using $\eta=\mathcal{O}(\frac{1}{T})$ and $b=\mathcal{O}(T)$ leads to an $\mathcal{O}(\frac{1}{\sqrt{T}})$ convergence rate, while using $\eta=\mathcal{O}(\frac{1}{T})$ and $b=\mathcal{O}(T^{2})$ leads to an $\mathcal{O}(\frac{1}{T})$ convergence rate. This implies that the larger $b$ is, the faster the convergence becomes. Such a trend can be seen in SGD and its variants (Smith et al., 2018; Umeda & Iiduka, 2025; Oowada & Iiduka, 2025). Motivated by this trend, we also consider using an increasing batch size, such as an exponentially growing batch size $b_{t}=b\delta^{t}$ (e.g., $\delta=2$ where the batch size is doubled every step (or epoch)). As a result, Muon has the upper bound $O(\frac{1}{T}+\eta)$, as seen in (R3) in Table 1 (Corollary 3.1(ii), (iv), and (vi)). Hence, using $\eta=\mathcal{O}(\frac{1}{T})$ leads to an $\mathcal{O}(\frac{1}{T})$ convergence rate. The convergence rate $\mathcal{O}(\frac{1}{T})$ of Muon shown in (R2) and (R3) is an improvement on the existing $\mathcal{O}(\frac{1}{\sqrt{T}})$ rate in Result (5) (Li & Hong, 2025, Theorem 2.1).
@@ 2.1 Nonconvex optimization problem and assumptions
We make the following standard assumptions to use an optimizer to minimize $f$ (see, e.g., (Umeda & Iiduka, 2025, Section 2.1) for justification of Assumption 2.1).
@@ 2.2 Muon optimizer
Let $b_{t}$ be the batch size at $t\in\mathbb{N}$ and let $\bm{\xi}_{t}\coloneqq(\xi_{t,1},\cdots,\xi_{t,i},\cdots,\xi_{t,b_{t}})^{\top}$ comprise $b_{t}$ independent and identically distributed variables and be independent of $\bm{W}\in\mathbb{R}^{m\times n}$. Assumption 2.1(ii) implies that the mini-batch stochastic gradient $\nabla f_{B_{t}}$ defined as in Step 2 of Algorithm 1 has the following properties for all $\bm{W}\in\mathbb{R}^{m\times n}$ and all $t\in\mathbb{N}$ (see, e.g., (Umeda & Iiduka, 2025, Proposition A.1)):
@@ 4 Conclusion
(iii) Section A.3 in (Umeda & Iiduka, 2025) ensures that $\eta_{t}=\frac{\eta}{2}(1+\cos\frac{t\pi}{T})$ satisfies
@@
(v) Section A.3 in (Umeda & Iiduka, 2025) ensures that $\eta_{t}=\eta(1-\frac{t}{T})^{p}$ satisfies

### v2

@@ 1.1 Background
Muon (Jordan et al., 2024) performs updates based on first-order information, similarly to stochastic gradient descent (SGD) (Robbins and Monro, 1951) and its momentum variants (Polyak, 1964; Nesterov, 1983; Sutskever et al., 2013), while distinguishing itself by incorporating an orthogonalization operation into the update direction (see (Si et al., 2025; Zhang et al., 2025; Huang et al., 2025; Liu et al., 2025b; Gruntkowska et al., 2025; Ahn et al., 2025) for the variants of Muon). Whereas Adam (Kingma and Ba, 2015) stabilizes training through coordinate-wise adaptive scaling, Muon projects the update matrix onto its orthogonal factor, thereby explicitly controlling the correlation structure of the update direction. This property enables Muon to preserve directional information even in high-dimensional spaces, achieving both numerical stability and efficient optimization. Empirical studies have further reported that Muon has advantages over Adam-type optimizers (Kingma and Ba, 2015; Loshchilov and Hutter, 2019) in terms of both computational efficiency and scalability (Liu et al., 2025a; Wang et al., 2025; AI et al., 2025; Mehta et al., 2025), highlighting its potential as a new optimizer capable of replacing conventional methods.
@@ 1.2 Motivation
Several attempts (Tang et al., 2025; Chang et al., 2025; Sato et al., 2025; Shen et al., 2025; Li and Hong, 2025) have been made to theoretically analyze the convergence properties of the Muon optimizer; the existing results are summarized in Table 1. These analyses aim to clarify the theoretical characteristics of Muon by establishing convergence guarantees or upper bounds under specific assumptions. However, each of these analyses is subject to limitations stemming from the strength of the imposed assumptions or the choice of evaluation criteria, and thus it does not necessarily fully capture the actual update structure of Muon.
Table 1: Comparison of the existing results (1)–(7) with our results (R1)(R5) for Muon. The existing results are (1) (Tang et al., 2025, Theorem 4.6), (2) (Chang et al., 2025, Theorem 3.4), (3) (Sato et al., 2025, Theorem 3.1), (4) (Shen et al., 2025, Theorem 4.1), (5) (Li and Hong, 2025, Theorem 2.1), (6) (Pethick et al., 2025a, Theorem 5.5), and (7) (Pethick et al., 2025b, Theorem 4.9), where Result (2) assumes the PL condition and Result (7) assumes $(L_{0},L_{1})$–smoothness condition. $(\bm{W}_{t})$ is the sequence generated by Muon and $\overline{\bm{W}}_{T}$ is the mean of $(\bm{W}_{t})_{t=0}^{T-1}$. $\|\cdot\|_{\mathrm{F}}$ is for the Frobenius norm and $\|\cdot\|_{*}$ is the dual norm of $\|\cdot\|_{\mathrm{F}}$. $\mathbb{E}$ is the total expectation. Let $\eta>0$, $b>0$, $\delta>1$, and $T$ be the number of steps. $O$ stands for Landau’s symbol, i.e., $y_{T}=O(x_{T})$ implies that, there exist $C>0$ and $t_{0}\in\mathbb{N}$ such that, for all $T\geq t_{0}$, $y_{T}\leq Cx_{T}$, while $y_{T}=\mathcal{O}(x_{T})$ implies that, for a fixed $T$, $y_{T}\leq Cx_{T}$. The notation $O(x(T,\eta,b))\hookrightarrow\mathcal{O}(x(T))$ implies that, when $\eta$ and $b$ are set by using the fixed $T$, the convergence rate changes from $O(x(T,\eta,b))$ to $\mathcal{O}(x(T))$.
@@
(6) $\mathbb{E}[\|\nabla f(\overline{\bm{W}}_{T})\|_{*}]$ $\frac{1}{2T^{3/4}}<\eta<\frac{1}{T^{3/4}}$ $b_{t}=b$ $\beta_{t}=\frac{1}{\sqrt{t}}$ $\mathcal{O}(\frac{1}{T^{1/4}})$
(7) $\mathbb{E}[\|\nabla f(\overline{\bm{W}}_{T})\|_{*}]$ $\eta=\mathcal{O}(\frac{1}{T^{3/4}})$ $b_{t}=b$ $\beta=\frac{1}{\sqrt{T}}$ $\mathcal{O}(\frac{1}{T^{1/4}})$
@@
Upon examining the results summarized in Table 1, we observe that Result (1) (Tang et al., 2025, Theorem 4.6) provides the convergence guarantee of Muon; however, its convergence rate would be inferior to that obtained in the present work. Result (2) (Chang et al., 2025, Theorem 3.4) can be interpreted as achieving a good convergence rate $O(\frac{(\log{T})^{2}}{T^{2/3}})$; however, it relies on the strong Polyak–Łojasiewicz (PL) condition, which significantly restricts its range of applicability. In the case of (3) (Sato et al., 2025, Theorem 3.1), a non-negligible variable $n$, where a parameter of a DNN is in $\mathbb{R}^{m\times n}$, remain in the analysis, and therefore a complete convergence result in a strict sense is not established. Result (4) (Shen et al., 2025, Theorem 4.1), similar to (1), provides a certain convergence guarantee, but its convergence rate would be again inferior to that achieved in this work. Result (5) (Li and Hong, 2025, Theorem 2.1) indicates that setting the number of steps $T$ as the batch size $b$ ensures a better $\mathcal{O}(\frac{1}{\sqrt{T}})$ convergence rate of Muon than an $\mathcal{O}(\frac{1}{T^{1/4}})$ convergence rate in Results (1) and (4). Result (6) (Pethick et al., 2025a, Theorem 5.5) and Result (7) (Pethick et al., 2025b, Theorem 4.9) considered algorithms similar to Muon and showed that the algorithms have an $\mathcal{O}(\frac{1}{T^{1/4}})$ convergence rate.
@@ 1.3 Contribution
Improved convergence rate for the Muon optimizer in the sense of the notation $\mathcal{O}$: We evaluate the upper bound in the sense of the notation $\mathcal{O}$, i.e., the case where the number of steps $T$ needed to train the DNN is fixed, by using practical learning rates and batch sizes. For example, when using a learning rate $\eta_{t}$ $(\leq\eta)$ defined by a constant, cosine-annealing, or polynomial decay learning rate and a constant batch size $b_{t}=b$, Muon has the upper bound $O(\frac{1}{T}+\eta+\frac{1}{\sqrt{b}})$, as shown in (R1) and (R2) in Table 1 (Corollary 3.1(i), (iii), and (v)). Hence, using $\eta=\mathcal{O}(\frac{1}{T})$ and $b=\mathcal{O}(T)$ leads to an $\mathcal{O}(\frac{1}{\sqrt{T}})$ convergence rate, while using $\eta=\mathcal{O}(\frac{1}{T})$ and $b=\mathcal{O}(T^{2})$ leads to an $\mathcal{O}(\frac{1}{T})$ convergence rate. This implies that the larger $b$ is, the faster the convergence becomes. Such a trend can be seen in SGD and its variants (Smith et al., 2018; Umeda and Iiduka, 2025; Oowada and Iiduka, 2025). Motivated by this trend, we also consider using an increasing batch size, such as an exponentially growing batch size $b_{t}=b\delta^{t}$ (e.g., $\delta=2$ where the batch size is doubled every step (or epoch)). As a result, Muon has the upper bound $O(\frac{1}{T}+\eta)$, as seen in (R3) in Table 1 (Corollary 3.1(ii), (iv), and (vi)). Hence, using $\eta=\mathcal{O}(\frac{1}{T})$ leads to an $\mathcal{O}(\frac{1}{T})$ convergence rate. The convergence rate $\mathcal{O}(\frac{1}{T})$ of Muon shown in (R2) and (R3) is an improvement on the existing $\mathcal{O}(\frac{1}{\sqrt{T}})$ rate in Result (5) (Li and Hong, 2025, Theorem 2.1) , Result (6) (Pethick et al., 2025a, Theorem 5.5), and Result (7) (Pethick et al., 2025b, Theorem 4.9).
@@ 2.1 Nonconvex optimization problem and assumptions
We make the following standard assumptions to use an optimizer to minimize $f$ (see, e.g., (Umeda and Iiduka, 2025, Section 2.1) for justification of Assumption 2.1).
@@ 2.2 Muon optimizer
Let $b_{t}$ be the batch size at $t\in\mathbb{N}$ and let $\bm{\xi}_{t}\coloneqq(\xi_{t,1},\cdots,\xi_{t,i},\cdots,\xi_{t,b_{t}})^{\top}$ comprise $b_{t}$ independent and identically distributed variables and be independent of $\bm{W}\in\mathbb{R}^{m\times n}$. Assumption 2.1(ii) implies that the mini-batch stochastic gradient $\nabla f_{B_{t}}$ defined as in Step 2 of Algorithm 1 has the following properties for all $\bm{W}\in\mathbb{R}^{m\times n}$ and all $t\in\mathbb{N}$ (see, e.g., (Umeda and Iiduka, 2025, Proposition A.1)):
@@ 4 Conclusion
(iii) Section A.3 in (Umeda and Iiduka, 2025) ensures that $\eta_{t}=\frac{\eta}{2}(1+\cos\frac{t\pi}{T})$ satisfies
@@
(v) Section A.3 in (Umeda and Iiduka, 2025) ensures that $\eta_{t}=\eta(1-\frac{t}{T})^{p}$ satisfies